# Maria Bazotte, Margarida Carvalho, and Thibaut Vidal

### v3

@@ 4.3 Standardization
A location-scale family of distributions has distributions with location $\mu$ and scale $\sigma$ parameters (Casella and Berger 2024, Chap. 3.5). The distributions of this family can be derived from a standard family distribution member with location and scale parameters equal to one and zero, respectively. If $\boldsymbol{\xi}^{u}$ follows a location-scale family distribution with parameters $\mu$ and $\sigma$ as first-stage decisions, then $t(\mu,\sigma,\boldsymbol{\vartheta}_{u})=\mu+\sigma\boldsymbol{\vartheta}_{u}$ is valid (Property 2 of Assumption 3). The $\boldsymbol{\vartheta}_{u}$ follows the standard family distribution member, which is independent of the values of $\mu$ and $\sigma$ (Property 1 of Assumption 3). If only the vector $h(\omega)$ has random elements, then Assumption 3 holds. Otherwise, a specific analysis of the terms $t_{1}(x,\widetilde{q}(\omega)^{\top})y$, $t_{3}(x,\widetilde{T}(\omega))x$ and $t_{4}(x,\widetilde{W}(\omega))y$ should be conducted. Many distributions have this property, such as the normal, generalized extreme value, exponential, and logistic distributions. Next, we exemplify this method for two of these distributions.
@@
If $\boldsymbol{\xi}^{u}\sim N(\mu,\sigma^{2})$ and $\mu$ and $\sigma$ are first-stage decisions, then we can apply the proposed transformation with $\boldsymbol{\vartheta}_{u}$ following the standard normal $N(0,1)$.
@@
If $\boldsymbol{\xi}^{u}\sim GEV(\mu,\sigma,\beta)$ with $\mu$ and $\sigma$ first-stage decisions and $\beta$ a fixed parameter, then we can apply the proposed transformation with $\boldsymbol{\vartheta}_{u}\sim GEV(0,1,\beta)$.
@@ 4.4 Convolution
Various random variables that follow well-known distributions can be expressed as the sum of other independent random variables. Notable examples include the binomial distribution, which is the sum of Bernoulli random variables; and the hypoexponential distribution, which is the sum of exponential random variables. Suppose that $\boldsymbol{\xi}_{x}^{u}$ is equal to the sum $\boldsymbol{\zeta}_{x}^{u1}+\boldsymbol{\zeta}_{x}^{u2}+\ldots+\boldsymbol{\zeta}_{x}^{un}$ of independent random variables $\boldsymbol{\zeta}_{x}^{uj}$’s following a specified endogenous distribution. In this case, the probability distribution of $\boldsymbol{\xi}_{x}^{u}$ is the convolution of the marginal distributions of the $\boldsymbol{\zeta}_{x}^{uj}$’s  (Hogg et al. 2013, Chap. 2). We can define a transformation $\boldsymbol{\zeta}_{x}^{uj}=\hat{t}_{j}(x,\boldsymbol{\vartheta}_{uj})$ for every endogenous random variable $\boldsymbol{\zeta}_{x}^{uj}$, $j\in\{1,\ldots,n\}$, independently. Thus, $t(x,\boldsymbol{\vartheta}_{u})=\sum_{j=1}^{n}\hat{t}_{j}(x,\boldsymbol{\vartheta}_{uj})$, with $\boldsymbol{\vartheta}_{u}=\{\boldsymbol{\vartheta}_{uj}\}_{j=1}^{n}$. Next, we demonstrate this transformation for the binomial distribution. We also provide an example of dependent random variables for which this methodology can be applied.
@@
We consider that $\boldsymbol{\xi}_{x}^{u}\sim Bin(n,\phi)$, where the probability of success of one trial $\phi\in[0,1]$ is a subvector of $x$, and the number of trials $n$ is a parameter. We have $\boldsymbol{\xi}_{x}^{u}=\boldsymbol{\zeta}_{x}^{u1}+\boldsymbol{\zeta}_{x}^{u2}+\ldots+\boldsymbol{\zeta}_{x}^{un}$, where the $\boldsymbol{\zeta}_{x}^{uj}$’s are endogenous and follow the $Bern(\phi)$ independently. We define the $\boldsymbol{\zeta}_{x}^{uj}$’s with the inversion for the Bernoulli distribution of Section 4.2.1. We have $\boldsymbol{\vartheta}_{uj}\sim U(0,1)$ independently for $j\in\{1,\ldots,n\}$. We set the variables $\upsilon_{ur}\in\{0,1\}$ equal to one if realization $\vartheta_{ur}<\phi$ and zero otherwise, and include Constraints (15) for all $\boldsymbol{\zeta}_{x}^{uj}$’s. Thus, $t(x,\vartheta_{u})=\sum_{j=1}^{n}\upsilon_{uj}$. The bilinear terms in $t_{1}(x,\widetilde{q}(\omega)^{\top})y$, $t_{3}(x,\widetilde{T}(\omega))x$, and $t_{4}(x,\widetilde{W}(\omega))y$ can be linearized by using the McCormick envelope (Assumption 3 holds). The inverse transformation (see Section 4.2.1) may be used in this case because the binomial probability of every realization is a convex function on the probability of success $\phi$ (i.e., we can apply Theorem 4.3). Both approaches have a similar number of variables and constraints and include big-M constraints. However, the general approach leads to convex (potentially nonlinear) constraints, whereas convolution leads to linear constraints.
Example 4.5
Consider two correlated random elements $\boldsymbol{\xi}_{x}^{A}$ and $\boldsymbol{\xi}_{x}^{B}$, such that $\boldsymbol{\xi}_{x}^{A}=\boldsymbol{\zeta}_{x}^{A1}+\boldsymbol{\zeta}_{x}^{AB2}$ and $\boldsymbol{\xi}_{x}^{B}=\boldsymbol{\zeta}_{x}^{B1}+\boldsymbol{\zeta}_{x}^{AB2}$. In this case, the correlation is a result of the dependency on $\boldsymbol{\zeta}_{x}^{AB2}$. Therefore, if $\boldsymbol{\zeta}_{x}^{AB2}=\hat{t}_{AB2}(x,\boldsymbol{\vartheta}_{AB2})$, $\boldsymbol{\zeta}_{x}^{A1}=\hat{t}_{A1}(x,\boldsymbol{\vartheta}_{A1})$, and $\boldsymbol{\zeta}_{x}^{B1}=\hat{t}_{B1}(x,\boldsymbol{\vartheta}_{B1})$, then $\boldsymbol{\xi}_{x}^{A}=\hat{t}(x,\boldsymbol{\vartheta}_{A1})+\hat{t}(x,\boldsymbol{\vartheta}_{AB2})$ and $\boldsymbol{\xi}_{x}^{B}=\hat{t}(x,\boldsymbol{\vartheta}_{B1})+\hat{t}(x,\boldsymbol{\vartheta}_{AB2})$, with $\boldsymbol{\vartheta_{A}}=\{\boldsymbol{\vartheta}_{A1},\boldsymbol{\vartheta}_{AB2}\}$ and $\boldsymbol{\vartheta_{B}}=\{\boldsymbol{\vartheta}_{B1},\boldsymbol{\vartheta}_{AB2}\}$.
@@ 4.5 Advantages and Disadvantages
The proposed method obtains the transformed Program (6), a two-stage stochastic program with exogenous uncertainty. This transformed program offers several advantages: (i) Solution procedures for such programs are extensively researched and can be applied based on their specific properties. (ii) Large-scale decomposition techniques, such as the L-shaped algorithm, are applicable depending on the characteristics of the transformed second-stage Program (6b). Besides, this method is versatile as various techniques can be employed to define the transformation function $t(x,\boldsymbol{\vartheta})$. As a result, multiple transformations may exist to model the same endogenous random vector. For instance, both the inverse and convolution methods can be applied to model the binomial random variable. The choice of a transformation should aim to achieve a transformed program with desirable properties. Furthermore, as demonstrated earlier in this section, in many cases, a transformation function leading to a mixed-integer linear or convex transformed Program (6) exists, even if the original Program (1) is nonlinear nonconvex. Even if the original program with endogenous uncertainty is an MILP, a simpler transformed program can still be obtained. For the discrete selection of distributions, the original program is an MILP obtained by linearizing the objective function with many variables and big-M constraints (see (Bhuiyan et al. 2020)). However, we propose a transformation based on the enumeration of distributions (see Section 4.1) that reduces considerably the need for additional variables or constraints.
Nevertheless, certain transformations, such as the inverse transformation for discrete random variables, may introduce many variables and constraints in the formulation. Hence, we explore the impact of such transformations in our experimental results. Additionally, the dimension and support of the exogenous random vector $\boldsymbol{\vartheta}$ are typically larger than those of the endogenous random vector $\boldsymbol{\xi}_{x}$. The convergence properties of sampling techniques depend on the support of $\boldsymbol{\vartheta}$, potentially requiring more samples. However, leveraging the same random samples of $\boldsymbol{\vartheta}$ for the SAA estimators $\hat{g}_{N}(x)$ for all $x\in\mathcal{X}$ to induce correlation between them is consistent with the efficiency-improvement technique of common random numbers (see discussion in Section 3.1).
@@ 5 Case Study
Pre-disaster or pre-interdiction problems with endogenous uncertainty have been investigated in the context of facility protection (Medal et al. 2016, Bhuiyan et al. 2020, Li et al. 2021, 2022) and network design protection (Viswanath et al. 2004, Flach 2010, Peeta et al. 2010, Du and Peeta 2014, Prestwich et al. 2014). To the best of our knowledge, these applications are the most recurrent in previous research. Thus, this section showcases our methodology by applying it to the NDFPP, which was introduced by Bhuiyan et al. (2020).
Bhuiyan et al. (2020) defined the endogenous random variables through a discrete selection of distributions. In contrast, we explore this and three other decision-dependent probability mass or density functions for the random variables. This section introduces our transformed program for the NDFPP, which has different versions according to the definition of the endogenous distribution.
@@ 5.1 Network Design and Facility Protection Problem
The NDFPP is a 2-SMIPD with complete recourse. First-stage decisions determine the allocation of protections for facilities and the network structure. Next, the scenario (natural disruptions) is realized, affecting the facilities’ capacity levels, and the second-stage decisions establish an optimal flow to satisfy client demands. The capacity levels of facilities are mutually independent endogenous random variables.
@@
We consider a set $\mathcal{D}_{0}$ of natural disruption events, which are different types of natural disasters with a center of occurrence. Facilities are exposed to distinct natural disruption intensity levels depending on their distance to the center of disruption $d\in\mathcal{D}_{0}$. The disruption intensity levels populate the set =$\mathcal{L}=\{1,\ldots,L\}$. We define the set $\mathcal{D}=\mathcal{D}_{0}\cup\{nd\}$ of possible events, where $nd$ is the no-natural-disruption event. The network is represented by the undirected graph $G=(\mathcal{N}_{\mathcal{T}},\mathcal{E})$, where $\mathcal{N}_{\mathcal{T}}=\mathcal{N_{F}}\cup\mathcal{N_{C}}\cup\{m\}$, with $\mathcal{N_{F}}$, $\mathcal{N_{C}}$, and $m$ being the set of facilities, the set of clients, and a dummy facility, respectively. In addition, $\mathcal{N}_{\mathcal{T}}^{0}=\mathcal{N}_{\mathcal{T}}\backslash\{m\}$ is the set of nodes without the dummy facility. We also specify a directed graph $G_{o}=(\mathcal{N}_{\mathcal{T}},\mathcal{A})$ by including the arcs $(i,j)$ and $(j,i)$ in $\mathcal{A}$ for each edge $[i,j]\in\mathcal{E}$. We denote as $e(i,j)$ the edge $e\in\mathcal{E}$ corresponding to arc $(i,j)\in\mathcal{A}$. Moreover, to represent the unsatisfied demand, we include arcs $(m,n)$ for every node $n\in\mathcal{N_{C}}$ in the set $\mathcal{A}$. Thus, the auxiliary set of arcs arriving at the dummy facility is defined as $\mathcal{A}_{m}=\{(m,n)\ \forall\ n\in\mathcal{N_{C}}\}$. The set $\mathcal{P}=\{1,\ldots,P\}$ contains the available protection levels for each facility, where larger values represent better protection. Finally, $\mathcal{S}$ is the set of scenarios, where each scenario is composed of an event and a realization of the exogenous random vector. We define the function $\hat{d}(s)$, which returns the event $d\in\mathcal{D}$ of scenario $s\in\mathcal{S}$. The realization of the exogenous random vector in scenario $s\in\mathcal{S}$ is denoted $\vartheta^{s}$. The definition of this random vector depends on the endogenous distribution and transformation function, so we define it later.
Protecting a facility $f\in\mathcal{N_{F}}$ with protection level $p\in\mathcal{P}$ has cost $c_{f}^{p}$ and opening edge $e\in\mathcal{E}$ has cost $c_{e}$. Every client node $n\in\mathcal{N_{C}}$ has demand $b_{n}$. We denote as $C$ the budget available for edge opening and facility protection. The per-unit-flow transportation cost of arc $(i,j)\in\mathcal{A}$ is defined as $q_{ij}$. The cost of arcs $\mathcal{A}_{m}$ leaving the dummy facility is $a\cdot\max_{(i,j)\in\mathcal{A}\backslash\mathcal{A}_{m}}q_{ij}$, where $a$ is the penalty per unit of unmet demand. Finally, $B=\sum_{n\in\mathcal{N_{C}}}b_{n}$ is the big-M value for edge capacity.
For the first stage, we define the binary variables $x_{f}^{p}$, which are equal to one if the protection level $p\in\mathcal{P}$ is allocated to facility $f\in\mathcal{N_{F}}$ and zero otherwise. We also include the binary variables $z_{e}$, equal to one if edge $e\in\mathcal{E}$ is installed and zero otherwise. For the second stage, we specify the variables $y_{ij}^{s}$ as the flow passing through arc $(i,j)\in\mathcal{A}$ in scenario $s\in\mathcal{S}$. The endogenous capacity of the facilities depends on the $x$ variables. Thus, following the notation of Section 4, we denote as $\xi_{x}^{fs}$ the endogenous capacity of facility $f\in\mathcal{N_{F}}$ in scenario $s\in\mathcal{S}$. We summarize this notation in Appendix 14. The general transformed program for the NDFPP is
$\min\mbox{: }$ $\mathbb{E}_{\boldsymbol{\vartheta}}[Q_{T}(x,\boldsymbol{\vartheta})]$ (16a)
@@
where the transformed second-stage program $Q_{T}(x,\vartheta_{s})$ for realization $\vartheta_{s}$ of $\boldsymbol{\vartheta}$ is
@@
In the first stage, the objective function (16a) minimizes the expected value of the transformed second stage program. We also have Constraints (16b) and (16c), which specify the selection of one level of protection for each facility and ensure that the cost budget for investing in facility protection and edge installation, respectively, is respected. The objective function (17a) of the second stage minimizes the average cost of transporting goods from facilities to clients. In addition, the second stage contains Constraints (17b)–(17e). These ensure flow conservation at client nodes (Constraints (17b)), the use only of available capacity (Constraints (17c)), the existence of flow only in installed arcs (Constraints (17d)), and a suitable definition of the endogenous realizations $\xi_{x}^{fs}$ (Constraints (17e)). We show the transformed SAA Program (7) of the NDFPP in Appendix 11.
In the following sections, we describe the different variants of this problem according to the considered decision-dependent distribution of $\xi_{x}^{fs}$. In other words, we detail Constraints (17e) and define the $\vartheta^{s}$. For the discrete endogenous distributions, we assume the post-disruption capacity of facilities to have $|\mathcal{W}|$ possible values, where $\mathcal{W}=\{0,\ldots,W\}$ is an ordered index set such that level 00 represents no capacity and level $W$ represents maximum capacity. The parameter $\nu_{w}$, with $w\in\mathcal{W}$, is the capacity at level $w$ such that $\nu_{w+1}=\nu_{w}+\overline{\nu}$; that is, $\overline{\nu}$ is the increase in capacity per unit increase in capacity level.
5.1.1 Discrete Selection of Distributions
First, we consider the discrete selection of distributions introduced by Bhuiyan et al. (2020). If a protection level $p\in\mathcal{P}$ is installed at facility $f\in\mathcal{N_{F}}$, then its capacity level is considered to follow $Bin(W,\overline{\phi}_{f}^{dp})$ when event $d\in\mathcal{D}$ is realized. The parameter $\overline{\phi}_{f}^{dp}$ is the probability of successfully increasing the capacity level by one. We follow the technique in Section 4.1. For facility $f\in\mathcal{N_{F}}$, protection $p\in\mathcal{P}$, and scenario $s\in\mathcal{S}$, we define $\vartheta^{s}_{fp}$ as the value of the realization of the capacity level, which follows the $Bin(W,\overline{\phi}_{f}^{dp})$. In this case, no extra variables are needed. Constraints (17e) are
$\xi_{x}^{fs}=\sum_{p\in\mathcal{P}}\vartheta^{s}_{fp}x_{f}^{p}$ $\forall\ f\in\mathcal{N_{F}}.$ (18)
Finally, this variant of the considered problem is denoted NDFPP-Selection.
5.1.2 Binomial Distribution

### v4

@@ 4.3 Standardization
A location-scale family of distributions has distributions with location $\mu$ and scale $\sigma$ parameters (Casella and Berger 2024, Chap. 3.5). The distributions of this family can be derived from a standard family distribution member with location and scale parameters equal to one and zero, respectively. If $\bm{\xi}^{u}$ follows a location-scale family distribution with parameters $\mu$ and $\sigma$ as first-stage decisions, then $t(\mu,\sigma,\bm{\vartheta}_{u})=\mu+\sigma\bm{\vartheta}_{u}$ is valid (Property 2 of Assumption 3). The $\bm{\vartheta}_{u}$ follows the standard family distribution member, which is independent of the values of $\mu$ and $\sigma$ (Property 1 of Assumption 3). If only the vector $h(\omega)$ has random elements, then Assumption 3 holds. Otherwise, a specific analysis of the terms $t_{1}(x,\widetilde{q}(\omega)^{\top})y$, $t_{3}(x,\widetilde{T}(\omega))x$ and $t_{4}(x,\widetilde{W}(\omega))y$ should be conducted. Thus, similar to the inversion, the transformed program may also be nonlinear and nonconvex, resulting in computational challenges. Many distributions belong to the location-scale family, including normal, generalized extreme value, exponential, and logistic distributions. Next, we exemplify this method for two of these distributions.
@@
If $\bm{\xi}^{u}\sim N(\mu,\sigma^{2})$ and $\mu$ and $\sigma$ are first-stage decisions, then we can apply the proposed transformation with $\bm{\vartheta}_{u}$ following the standard normal $N(0,1)$.
@@
If $\bm{\xi}^{u}\sim GEV(\mu,\sigma,\beta)$ with $\mu$ and $\sigma$ first-stage decisions and $\beta$ a fixed parameter, then we can apply the proposed transformation with $\bm{\vartheta}_{u}\sim GEV(0,1,\beta)$.
@@ 4.4 Convolution
Various random variables that follow well-known distributions can be expressed as the sum of other independent random variables. Notable examples include the binomial distribution, which is the sum of Bernoulli random variables, and the hypoexponential distribution, which is the sum of exponential random variables. Suppose that $\bm{\xi}_{x}^{u}$ is equal to the sum $\bm{\zeta}_{x}^{u1}+\bm{\zeta}_{x}^{u2}+\ldots+\bm{\zeta}_{x}^{un}$ of independent random variables $\bm{\zeta}_{x}^{uj}$’s following a specified endogenous distribution. In this case, the probability distribution of $\bm{\xi}_{x}^{u}$ is the convolution of the marginal distributions of the $\bm{\zeta}_{x}^{uj}$’s (Hogg et al. 2013, Chap. 2). We can define a transformation $\bm{\zeta}_{x}^{uj}=\hat{t}_{j}(x,\bm{\vartheta}_{uj})$ for every endogenous random variable $\bm{\zeta}_{x}^{uj}$, $j\in\{1,\ldots,n\}$, independently. Thus, $t(x,\bm{\vartheta}_{u})=\sum_{j=1}^{n}\hat{t}_{j}(x,\bm{\vartheta}_{uj})$, with $\bm{\vartheta}_{u}=\{\bm{\vartheta}_{uj}\}_{j=1}^{n}$. The complexity of the problem depends on the definition of each transformation $\hat{t}_{j}(x,\bm{\vartheta}_{uj})$ and the number of components $n$, thus limiting the application of this method by the size of $n$. Next, we demonstrate this transformation for the binomial distribution. We also provide an example of dependent random variables for which this methodology can be applied in Appendix 9 (Example 9.4).
@@
We consider that $\bm{\xi}_{x}^{u}\sim Bin(n,\phi)$, where the probability of success of one trial $\phi\in[0,1]$ is a subvector of $x$, and the number of trials $n$ is a parameter. We have $\bm{\xi}_{x}^{u}=\bm{\zeta}_{x}^{u1}+\bm{\zeta}_{x}^{u2}+\ldots+\bm{\zeta}_{x}^{un}$, where the $\bm{\zeta}_{x}^{uj}$’s are endogenous and follow the $Bern(\phi)$ independently. We define the $\bm{\zeta}_{x}^{uj}$’s with the inversion for the Bernoulli distribution of Section 4.1.1. We have ${\bm{\vartheta}_{uj}\sim U(0,1)}$ independently for $j\in\{1,\ldots,n\}$. We set the variables $\upsilon_{ur}\in\{0,1\}$ equal to one if realization $\vartheta_{ur}<\phi$ and zero otherwise, and include Constraints (14) for all $\bm{\zeta}_{x}^{uj}$’s. Thus, $t(x,\vartheta_{u})=\sum_{j=1}^{n}\upsilon_{uj}$. If only $h(\omega)$ is uncertain, then Assumption 3 is respected. Otherwise, the bilinear terms in $t_{1}(x,\widetilde{q}(\omega)^{\top})y$, $t_{3}(x,\widetilde{T}(\omega))x$, and $t_{4}(x,\widetilde{W}(\omega))y$ can be linearized by using the McCormick envelope (Assumption 3 holds). In this case, the application of this technique is limited to the dimensions of $y$ and $x$ due to the increase in the number of variables. Furthermore, inversion (see Section 4.1.1) may be used in this case because the binomial probability of every realization is a convex function on the probability of success $\phi$ (i.e., we can apply Theorem 4.1). Both approaches have a similar number of variables and constraints and include big-M constraints. However, inversion leads to convex (potentially nonlinear) constraints, whereas convolution leads to linear constraints.
@@ 4.5 Advantages and Disadvantages
The transformed Program (4), a two-stage program with exogenous uncertainty, offers several advantages: (i) Solution procedures for such programs are extensively researched and can be applied based on their specific properties. (ii) Large-scale decomposition techniques, such as the L-shaped algorithm, are applicable depending on the characteristics of the transformed second-stage Program (5). (iii) Sampling techniques become straightforward, enabling the treatment of problems with continuous endogenous random variables or discrete ones with many realizations. In contrast, using sampling to solve the original program with endogenous uncertainty is challenging, as the distribution depends on first-stage decisions. Besides, our method is versatile as various techniques can be employed to define the transformation function $t(x,\bm{\vartheta})$. As a result, multiple transformations may exist to model the same endogenous random vector. For instance, both the inverse and convolution methods can be applied to model the binomial random variable. The choice of a transformation should aim to achieve a transformed program with desirable properties. Furthermore, as demonstrated earlier in this section, in many cases, a transformation function leading to a mixed-integer linear or convex transformed Program (4) exists, even if the original Program (1) is nonlinear nonconvex. This is particularly the case for problems where only the vector $h(\omega)$ is uncertain, as the transformation $t_{2}\bm{(}x,\widetilde{h}\left(\omega\right)\bm{)}$ does not multiply the variables $x$ or $y$ in Program (5). For the discrete selection of distributions where only $h(\omega)$ is uncertain, the original program is an MILP obtained by linearizing the objective function with many variables and big-M constraints (see (Bhuiyan et al. 2020)). However, we propose a transformation based on the enumeration of distributions that reduces the need for additional variables or constraints in this case.
Nevertheless, certain transformations, such as enumeration and the inverse transformation for discrete random variables, may introduce in the formulation many constraints and variables, which are often integer when dealing with discrete distributions. Hence, we explore the impact of such transformations in our experimental results. For problems where the cost vector $q(\omega)$, technology matrix $T(\omega)$, and/or recourse matrix $W(\omega)$ are uncertain, obtaining a transformed program with a convex relaxation is more challenging due to the terms $t_{1}(x,\widetilde{q}(\omega)^{\top})y$, $t_{3}(x,\widetilde{T}(\omega))x$, and $t_{4}(x,\widetilde{W}(\omega))y$. For discrete distributions, introducing integer variables to define the transformations enables the use of McCormick linearization as a last resort. However, this approach is constrained by the dimensions of $x$, $y$, and the number of uncertain parameters. Additionally, the dimension and support of the exogenous random vector $\bm{\vartheta}$ are typically larger than those of the endogenous random vector $\bm{\xi}_{x}$. The convergence properties of sampling techniques depend on the support of $\bm{\vartheta}$, potentially requiring more samples. However, leveraging the same random samples of $\bm{\vartheta}$ for the SAA estimators $\hat{g}_{N}(x)$ for all $x\in\mathcal{X}$ to induce correlation between them is consistent with the efficiency-improvement technique of common random numbers (see discussion in Section 3.1).
@@ 5 Case Study
Pre-disaster or pre-interdiction problems with endogenous uncertainty have been investigated in the context of facility protection (Medal et al. 2016, Bhuiyan et al. 2020, Li et al. 2021, 2022) and network design protection (Viswanath et al. 2004, Flach 2010, Peeta et al. 2010, Du and Peeta 2014, Prestwich et al. 2014). To the best of our knowledge, these applications are the most recurrent in previous research. Thus, this section showcases our methodology by applying it to the NDFPP, which was introduced by Bhuiyan et al. (2020)Bhuiyan et al. (2020) defined the endogenous random variables through a discrete selection of distributions. In contrast, we explore this and three other decision-dependent probability mass or density functions for the random variables. This section introduces our transformed program for the NDFPP, which has different versions according to the definition of the endogenous distribution.
@@ 5.1 Network Design and Facility Protection Problem
The NDFPP is a 2-SMIPD with complete recourse. First-stage decisions determine facility protection allocations and the network structure. Next, the scenario (natural disruptions) is realized, affecting facility capacity levels, and second-stage decisions establish an optimal flow to satisfy client demands. Facility capacity levels are mutually independent endogenous random variables.
@@
We consider a set $\mathcal{D}_{0}$ of natural disruption events, representing different types of natural disasters centered at specific locations. Facilities experience varying disruption intensities based on their distance from the disruption center $d\in\mathcal{D}_{0}$, with intensity levels in the set =$\mathcal{L}=\{1,\ldots,L\}$. We define the set of possible events $\mathcal{D}=\mathcal{D}_{0}\cup\{nd\}$, which includes the no-natural-disruption event $nd$. The network is represented by the undirected graph $G=(\mathcal{N}_{\mathcal{T}},\mathcal{E})$, where $\mathcal{N}_{\mathcal{T}}=\mathcal{N_{F}}\cup\mathcal{N_{C}}\cup\{m\}$ consists of facilities ($\mathcal{N_{F}}$), clients ($\mathcal{N_{C}}$), and a dummy facility ($m$). The set $\mathcal{N}_{\mathcal{T}}^{0}=\mathcal{N}_{\mathcal{T}}\backslash\{m\}$ excludes the dummy facility. We also specify a directed graph $G_{o}=(\mathcal{N}_{\mathcal{T}},\mathcal{A})$, including arcs $(i,j)$ and $(j,i)$ for each edge $[i,j]\in\mathcal{E}$. The edge corresponding to arc $(i,j)\in\mathcal{A}$ is denoted as $e(i,j)$. To represent unsatisfied demand, we include arcs $(m,n)$ for every client node $n\in\mathcal{N_{C}}$ in $\mathcal{A}$, with the auxiliary set of arcs arriving at the dummy facility defined as $\mathcal{A}_{m}=\{(m,n)\ \forall\ n\in\mathcal{N_{C}}\}$. The set $\mathcal{P}=\{1,\ldots,P\}$ contains available protection levels for each facility, where larger values indicate better protection. Finally, $\mathcal{S}$ is the set of scenarios, each composed of an event and a realization of the exogenous random vector. The function $\hat{d}(s)$ returns the event $d\in\mathcal{D}$ for scenario $s\in\mathcal{S}$. The realization of the exogenous random vector in scenario $s$ is denoted $\vartheta^{s}$. The definition of this random vector depends on the endogenous distribution and transformation function, which we define later. Moreover, protecting a facility $f\in\mathcal{N_{F}}$ with protection level $p\in\mathcal{P}$ costs $c_{f}^{p}$, and opening an edge $e\in\mathcal{E}$ costs $c_{e}$. Each client node $n\in\mathcal{N_{C}}$ has a demand $b_{n}$. The budget for edge opening and facility protection is denoted by $C$. The per-unit-flow transportation cost for arc $(i,j)\in\mathcal{A}$ is $q_{ij}$. Moreover, the cost for arcs $\mathcal{A}_{m}$ leaving the dummy facility is $a\cdot\max_{(i,j)\in\mathcal{A}\backslash\mathcal{A}_{m}}q_{ij}$, where $a$ is the penalty per unit of unmet demand. The big-M value for edge capacity is $B=\sum_{n\in\mathcal{N_{C}}}b_{n}$. For the first stage, we define the binary variables $x_{f}^{p}$, which are equal to one if protection level $p\in\mathcal{P}$ is allocated to facility $f\in\mathcal{N_{F}}$ and zero otherwise. We also include the binary variables $z_{e}$, which are equal to one if edge $e\in\mathcal{E}$ is installed and zero otherwise. For the second stage, we specify the variables $y_{ij}^{s}$, which represent the flow through arc $(i,j)\in\mathcal{A}$ in scenario $s\in\mathcal{S}$. The endogenous capacity of facilities, $\xi_{x}^{fs}$ for facility $f\in\mathcal{N_{F}}$ in scenario $s\in\mathcal{S}$, depends on the $x$ variables. We summarize this notation in Appendix 15. The transformed program for the NDFPP is
$\min\mbox{: }$ $\mathbb{E}_{\bm{\vartheta}}[Q_{T}(x,\bm{\vartheta})]$ (15a)
@@
where the transformed second-stage program $Q_{T}(x,\vartheta_{s})$ for realization $\vartheta_{s}$ of $\bm{\vartheta}$ is
@@
In the first stage, the objective function (15a) minimizes the expected value of the transformed second-stage program. We also have Constraints (15b) and (15c), which specify the selection of one level of protection for each facility and ensure that the cost budget for investing in facility protection and edge installation, respectively, is respected. The objective function (16a) of the second stage minimizes the average cost of transporting goods from facilities to clients. In addition, the second stage contains Constraints (16b)–(16e). These ensure flow conservation at client nodes (Constraints (16b)), the use only of available capacity (Constraints (16c)), the existence of flow only in installed arcs (Constraints (16d)), and a suitable definition of the endogenous realizations $\xi_{x}^{fs}$ (Constraints (16e)). We note that, following the notation in Program (1), only $h(\omega)$ is uncertain in this problem. We show the transformed SAA Program (6) of the NDFPP in Appendix 11.
In the following sections, we describe the various problem variants based on the decision-dependent distribution of $\xi_{x}^{fs}$. Specifically, we detail Constraints (16e) and define $\vartheta^{s}$. For discrete endogenous distributions, we assume that the post-disruption capacity of facilities can take on $|\mathcal{W}|$ possible values, where $\mathcal{W}=\{0,\ldots,W\}$ is an ordered index set. Here, level 0 represents no capacity, and level $W$ represents maximum capacity. The parameter $\nu_{w}$ for $w\in\mathcal{W}$ denotes the capacity at level $w$, with $\nu_{w+1}=\nu_{w}+\overline{\nu}$, where $\overline{\nu}$ is the increase in capacity per unit increase in capacity level.
5.1.1 Discrete Selection of Distributions.
First, we consider the discrete selection of distributions introduced by Bhuiyan et al. (2020). If a protection level $p\in\mathcal{P}$ is installed at facility $f\in\mathcal{N_{F}}$, then its capacity level is considered to follow $Bin(W,\overline{\phi}_{f}^{dp})$ when event $d\in\mathcal{D}$ is realized. The parameter $\overline{\phi}_{f}^{dp}$ is the probability of successfully increasing the capacity level by one. We follow the enumration in Section 4.2. For facility $f\in\mathcal{N_{F}}$, protection $p\in\mathcal{P}$, and scenario $s\in\mathcal{S}$, we define $\vartheta^{s}_{fp}$ as the value of the realization of the capacity level following the $Bin(W,\overline{\phi}_{f}^{dp})$. In this case, no extra variables are needed. We have $\xi_{x}^{fs}=\sum_{p\in\mathcal{P}}\vartheta^{s}_{fp}x_{f}^{p}$ for every facility $f\in\mathcal{N_{F}}$, which model Constraints (16e). Finally, this variant of the considered problem is denoted NDFPP-Selection.
5.1.2 Binomial Distribution.